A student washes her clothes at a laundromat once a week. The five number summary for the time (in minutes) she spent in the laundromat for each of 15 randomly selected weeks was 62, 73, 83, 93, and 115.

What's the lower fence?

a) 105

b) 43

c) 65

d) 123

e) 93
Answer #1

Five number summery = min , Q1 , Q2(median) , Q3 , Max.

First quartile Q1 = 73

Third quartile Q3 = 93

IQR = Q3 - Q1 = 93 - 73 = 20

Lower fence = Q1 - 1.5 * IQR

= 73 - 1.5 * 20

= 43